# Locally Repairable Convertible Codes: Improved Lower Bound and General Construction

### v1

@@ IV A General Construction of LRCC
To obtain the parity check matrix of the final code, we concatenate the parity check matrices $\overline{H}_{1},\dots,\overline{H}_{t}$ and derive the following matrix:
$H_{F}=\begin{pmatrix}\bm{W}_{11}&\bm{O}&\cdots&\bm{O}&\bm{O}\\ \bm{O}&\bm{W}_{21}&\cdots&\bm{O}&\bm{O}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \bm{O}&\bm{O}&\cdots&\bm{W}_{t1}&\bm{O}\\ \bm{O}&\bm{O}&\cdots&\bm{O}&\bm{W}_{12}\\ \bm{N}_{11}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{11}&\bm{N}_{21}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{21}&\cdots&\bm{N}_{t1}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{t1}&\bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}.$ (17)
@@
Now, we construct the final code $\mathcal{C}_{F}$. Choose any $t$ codewords $\bm{c}_{1}\in\mathcal{C}_{1},\dots,\bm{c}_{t}\in\mathcal{C}_{t}$, then we can calculate the new symbols as
@@
and the code symbols $\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]}$ corresponds to the read symbols for $\bm{c}_{i}$ for $1\leqslant i\leqslant t$. The final code $\mathcal{C}_{F}$ is defined as
@@
Let the code $\tilde{\mathcal{C}}$ be given by the parity check matrix $H_{F}$. We first show that $\mathcal{C}_{F}\subseteq\tilde{\mathcal{C}}$. On the one hand, according to (19) and the unchanged symbols satisfy the local parity check equations, for any codeword $\bm{c}\in\mathcal{C}_{F}$ and $i\in[\ell]$, we have
$(\underset{i-1\text{ times}}{\underbrace{O,\cdots,O}},A_{i},\underset{\ell+h-i\text{ times}}{\underbrace{O,\cdots,O}})\cdot\bm{c}^{\top}=\bm{0}.$
@@
$(\underset{i-1\text{ times}}{\underbrace{\bm{O},\cdots,\bm{O}}},\bm{W}_{i1},\underset{t-i+1\text{ times}}{\underbrace{\bm{O},\cdots,\bm{O}}})\cdot\bm{c}^{\top}=\bm{0}.$ (20)
@@
$\begin{split}&\begin{pmatrix}\bm{O}&\cdots&\bm{O}&\bm{W}_{12}\\ \bm{N}_{11}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{11}&\cdots&\bm{N}_{t1}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{t1}&\bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}\cdot\bm{c}^{\top}\\ &=\sum_{i=1}^{t}\begin{pmatrix}\bm{O}\\ \bm{N}_{i1}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{i1}\end{pmatrix}\cdot(\bm{c}_{i}|_{[s(r+\delta-1)]})^{\top}+\begin{pmatrix}\bm{W}_{12}\\ \bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}\cdot(\bm{c}^{*})^{\top}\\ &=-\sum_{i=1}^{t}\begin{pmatrix}\bm{W}_{12}\\ \bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}\cdot(\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]})^{\top}+\begin{pmatrix}\bm{W}_{12}\\ \bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}\cdot(\bm{c}^{*})^{\top}\\ &=\begin{pmatrix}\bm{W}_{12}\\ \bm{N}_{12}-\bm{N}_{13}\bm{M}_{13}^{-1}\bm{M}_{12}\end{pmatrix}\cdot\left(\bm{c}^{*}-\sum_{i=1}^{t}\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]}\right)^{\top}\\ &=\bm{0},\end{split}$ (21)
where the second equation holds by $\overline{H}_{i}\cdot\bm{c}_{i}|_{[(s+h)(r+\delta-1)]}^{\top}=\bm{0}$ for $i\in[t]$, and the last equation comes from (18). Combining (20) and (21), we have that $H_{F}\cdot\bm{c}^{\top}=\bm{0}$ for any $\bm{c}\in\mathcal{C}_{F}$. Therefore, $\mathcal{C}_{F}\subseteq\tilde{\mathcal{C}}$ and
$\ell r=\dim(\mathcal{C}_{F})\leqslant\dim(\tilde{\mathcal{C}}).$ (22)
Secondly, we claim that $\mathcal{C}_{F}$ is an optimal $[(\ell+h)(r+\delta-1),\ell r,d]_{q}$-LRC. By the fact $\mathcal{C}_{F}\subseteq\tilde{\mathcal{C}}$, it is sufficient to show that $\mathcal{C}_{F}=\tilde{\mathcal{C}}$ and $\tilde{\mathcal{C}}$ is an optimal LRC. Note that $H_{F}$ is also the parity check matrix of the puncturing code $\mathcal{C}|_{[(\ell+h)(r+\delta-1)]}$ from $\mathcal{C}$ by (16) and (17). Thus, $\tilde{\mathcal{C}}=\mathcal{C}|_{[(\ell+h)(r+\delta-1)]}$. Recall that $\mathcal{C}$ is an optimal LRC, which means
@@
On the other hand, $\tilde{\mathcal{C}}$ is an $[(\ell+h)(r+\delta-1),\ell r,d(\tilde{\mathcal{C}})]_{a}$-LRC with $(r,\delta)_{a}$-locality, by the Singleton-type bound (3), we have
@@
Combining (18) and (19), we derive that the write access cost is $h(r+\delta-1)$ and the read access cost is $thr$. According to Theorem 2, we have that
$\rho_{w}$ $\geqslant t(\delta+h(r+\delta-1)+(t-1)sr-1+(s(t-1)-1)(\delta-1))$
@@
$\Delta$ $=(\ell+h)(r+\delta-1)-2(\delta+h(r+\delta-1))$
@@
$\rho_{r}$ $\geqslant t\left(sr-\Delta+(\delta-1)\lfloor\frac{\Delta}{r+\delta-1}\rfloor\right)$
$=t(sr+h(r+\delta-1)-s(r+\delta-1)-\delta+1+(\delta-1)(s-h-1))$
@@
The read and write access costs both achieve the lower bound in Theorem 2. Thus, the LRCC above is access-optimal, which completes the proof. ∎
@@
The $t$ initial codes in Theorem 3 can be the same one code, when $h=|\mathcal{P}|=g-\ell$ and $H||_{\mathcal{P}_{i}}=D_{i}\cdot H||_{\mathcal{P}_{1}}$ for $i\in[t]$ where $D_{i}$ is an invertible diagonal matrix as
$D_{i}=\begin{pmatrix}D_{i1}&O_{s(\delta-1)\times hr}\\ O_{hr\times s(\delta-1)}&D_{i2}\end{pmatrix}.$
In this case, both $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ and $A_{i}\cdot H||_{\mathcal{P}_{1}\cup\mathcal{P}}$ can be the parity check matrices of initial code $\mathcal{C}_{i}$, where
$A_{i}=\begin{pmatrix}D_{i1}&O_{s(\delta-1)\times h(\delta-1)}&O_{s(\delta-1)\times hr}\\ O_{h(\delta-1)\times s(\delta-1)}&I_{h(\delta-1)\times h(\delta-1)}&O_{h(\delta-1)\times hr}\\ O_{hr\times s(\delta-1)}&O_{hr\times h(\delta-1)}&D_{i2}\end{pmatrix}.$
Therefore, for ever $1\leqslant i\leqslant t$,
$\overline{H}_{i}=H||_{\mathcal{P}_{i}\cup\mathcal{P}}=\begin{pmatrix}\bm{W}_{i1}&O\\ O&\bm{W}_{12}\\ \bm{N}_{i1}&\bm{N}_{12}\end{pmatrix}.$
And the conversion procedure function (18) to calculate the new symbols will becomes
$\bm{c}^{*}=\begin{pmatrix}\bm{W}_{12}\\ \bm{N}_{12}\end{pmatrix}^{-1}\sum_{i=1}^{t}\left(\begin{pmatrix}I_{h(\delta-1)\times h(\delta-1)}&O_{h(\delta-1)\times hr}\\ O_{hr\times s(\delta-1)}&D_{i2}\end{pmatrix}\begin{pmatrix}\bm{W}_{12}\\ \bm{N}_{12}\end{pmatrix}\cdot(\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]})^{\top}\right).$
Also the equation (20) and (21) hold in this case. Therefore, the new LRCC is access-optimal, and the initial and final codes are optimal LRCs.
V Explicitly Constructions of LRCCs
Recall that the proof of Theorem 3 gives the initial codes by means of the parity check matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ as (15) for $i\in[t]$, the final code by means of $H_{F}$ as (17), and the conversion procedure as (18). Also, the initial and final codes are all optimal LRCs, and the derived LRCC is access-optimal. In this section, we show two explicit constructions and it is sufficient to show the initial and final codes are all with super-linear length or are all maximally recoverable codes.
@@ V-A Super-linear length
For MDS codes, the MDS conjecture implies that the length of MDS codes may not exceed $q+1$ ($q+2$ for $2\mid q$ and $k=3$), which means for MDS convertible codes the length of final codes may be upper bounded by $q+1$ ($q+2$ for $2\mid q$ and $k=3$). Analogously, Kong in [23] raises an interesting question about how small the field size can be to ensure the existence of LRCCs with optimal access costs.
The trivial answer for this question is the length of final codes for LRCCs can not exceed the optimal length of LRC referring the known bounds in [5, 24].
Lemma 8 ([5, Theorem 2]):
Let $n=w(r+\delta-1)$, $\delta\geqslant 2$, $k=ur+v$, and additionally, $r|k$ or $u\geqslant 2(r+1-v)$, where all parameters are integers. Assume there exists an optimal $[n,k,d]_{q}$ linear code $\mathcal{C}$ with all symbol $(r,\delta)$-locality, and define $t=\left\lfloor(d-1)/\delta\right\rfloor$. If $42t+1>$ 4, then
$n$$\leqslant\begin{cases}\frac{r+\delta-1}{r}\left(\frac{t-1}{2(q-1)}q^{\frac{2(w-u)r-2v-2}{t-1}}+1\right),&\text{ if }t\text{ is odd},\\ \frac{t(r+\delta-1)}{2r(q-1)}q^{\frac{2(w-u)r-2v}{t}},&\text{ if }t\text{ is even},\\ \end{cases}$
where $w-u$ can also be rewritten as $w-u=\lfloor(d-1+v)/(r+\delta-1)\rfloor$.
In this subsection, we propose an explicit construction, which implies that the above bound is also tight asymptotically for LRCCs for some special cases.
Base Code A ([24, Construction A]): Let $d\geqslant\delta+1$ and $n=m(r+\delta-1)$, where $d,\delta,r,m$ are positive integers. For $i\in[m]$, let $G_{i}=\{g_{i,1},g_{i,2},\dots,g_{i,r+\delta-1}\}$ be an $(r+\delta-1)$-subset of $\mathbb{F}_{q}$. Then, for each $i\in[m]$, we can construct a $(d-1)\times(r+\delta-1)$ Vandermonde matrix with generating set $G_{i}$ of the form $(U_{i}^{\top},V_{i}^{\top})^{\top}$, where
@@
$H=\begin{pmatrix}U_{1}&O&\cdots&O\\ O&U_{2}&\cdots&O\\ \vdots&\vdots&\ddots&\vdots\\ O&O&\cdots&U_{m}\\ V_{1}&V_{2}&\cdots&V_{m}\end{pmatrix}.$ (25)
@@
Lemma 9 ([24, Theorem III.4]):
Let $r+\delta\geqslant d\geqslant 2\delta+1$. Suppose that for any subset $S\subseteq[m]$ with $2\leqslant|S|\leqslant\lfloor\frac{d-1}{\delta}\rfloor$, we have
@@
then any $d-1$ columns of $H$ given by (25) are linear independent. As a result, the base code $\mathcal{C}$ is an optimal $[n,k,d]_{q}$-LRC with $(r,\delta)_{a}$-locality.
@@
Subsequent improved Singleton-type bound is employed to determine the dimension of the code.
Lemma 10 ([24, Remark II.4]):
$d\leqslant n-k+1-\left\lceil\frac{k}{r}\right\rceil(\delta-1).$
Therefore, we can derive the following corollary.
Corollary 1:
Let $m\geqslant v\geqslant 1$, and $i_{1}<i_{2}<\cdots<i_{v}$ where $i_{j}\in[m]$ for $j\in[v]$. Then the code $\mathcal{C}$ given by the parity check matrix $H||_{\{i_{1},i_{2},\dots,i_{v}\}}$ is an optimal $[v(r+\delta-1),(v-1)r,r+\delta]_{q}$-LRC with $(r,\delta)_{a}$-locality, where $H$ is the parity check matrix of the code $\tilde{\mathcal{C}}$ from Base Code A satisfying $r+\delta=d\geqslant 2\delta+1$ and condition (26).
@@
Since $G_{i},i\in[m]$ satisfies condition (26), then $G_{j},j\in{\{i_{1},i_{2},\dots,i_{v}\}}\subseteq[m]$ satisfies condition (26). Also, the $\overline{d}$ for $\mathcal{C}$ satisfies $\overline{d}=d=r+\delta\geqslant 2\delta+1$. According to Lemma 9 and Lemma 10, the dimension $k(\mathcal{C})$ of $\mathcal{C}$ with parameters $[v(r+\delta-1),k(\mathcal{C}),\overline{d}=r+\delta]_{q}$ satisfies
@@
Corollary 2:
Let $u$ be a positive integer. Let the code $\mathcal{C}$ be from Base Code A satisfying $r+\delta=d\geqslant 2\delta+1$ and condition (26). If $u\mid m-1$, then there exists an access-optimal LRCC with parameters
@@
Moreover, the initial and final code are with super-linear length.
@@
Let $\mathcal{P}_{i}=\{(i-1)u+1,\dots,iu\}$ for $i\in[\frac{m-1}{u}]$, and $\mathcal{P}=\{m\}$. Then $\mathcal{C}_{i}$ given by the parity check matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ for $i\in[\frac{m-1}{u}]$ are the initial codes, where $H$ is given by (25). According to the proof of Theorem 3, the conversion procedure is (18) and the final code is $\mathcal{C}$. Therefore, there exists an access-optimal LRCC. According to Corollary 1, the initial codes are optimal $[(u+1)(r+\delta-1),ur,r+\delta]_{q}$-LRC with $(r,\delta)_{a}$-locality.
It is well-known that the base code $\mathcal{C}$ can be with super-linear length [24]. Therefore, the code length $m(r+\delta-1)$ satisfies that
@@
Since $(u+1)(r+\delta-1)=\frac{u+1}{m}\cdot m(r+\delta-1)$ and $\frac{u+1}{m}$ is a constant. Therefore, the initial codes can be with super-linear length. ∎
@@
For the case $2\delta+1=d$, $r=\delta+1$, the parameters in Lemma 8 can be figure out as $w-u=u+1-u=1$, $t=2$ and
@@
Thus, the code length $\Omega(q^{\delta})$ is asymptotically optimal with respect to the bound in Lemma 8.
@@
When arbitrary two of $\cup_{j\in\mathcal{P}_{i}}G_{j}$ for $i\in[t]$ are belong to different cosets of one multiplicative group of $\mathbb{F}_{q}^{*}$ in Corollary 2, i.e.,

### v2

@@ IV A General Construction of LRCCs
Construct the final code: To obtain the parity check matrix of the final code, we concatenate the parity check matrices $\overline{H}_{1},\dots,\overline{H}_{t}$ and derive the following matrix:
$H_{F}=\begin{pmatrix}W_{11}&\bm{0}&\cdots&\bm{0}&\bm{0}\\ \bm{0}&W_{21}&\cdots&\bm{0}&\bm{0}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \bm{0}&\bm{0}&\cdots&W_{t1}&\bm{0}\\ \bm{0}&\bm{0}&\cdots&\bm{0}&W_{12}\\ N_{11}-N_{13}M_{13}^{-1}M_{11}&N_{21}-N_{13}M_{13}^{-1}M_{21}&\cdots&N_{t1}-N_{13}M_{13}^{-1}M_{t1}&N_{12}-N_{13}M_{13}^{-1}M_{12}\end{pmatrix}$. (27)
@@
Now, we construct the final code $\mathcal{C}_{F}$. Choose any $t$ codewords $\bm{c}_{1}\in\mathcal{C}_{1},\dots,\bm{c}_{t}\in\mathcal{C}_{t}$, and calculate the new symbols as
@@
and the code symbols $\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]}$ correspond to the read symbols for $\bm{c}_{i}$ for $1\leqslant i\leqslant t$. The final code $\mathcal{C}_{F}$ is defined as
@@
Verify that the final code is an optimal LRC: Let the code $\tilde{\mathcal{C}}$ be given by the parity check matrix $H_{F}$. We first show that $\mathcal{C}_{F}\subseteq\tilde{\mathcal{C}}$. On the one hand, according to (29) and the unchanged symbols satisfy the local parity check equations, for any codeword $\bm{c}\in\mathcal{C}_{F}$ and $i\in[\ell]$, we have
$(\underset{i-1\text{ times}}{\underbrace{\bm{0},\cdots,\bm{0}}},A_{i},\underset{\ell+h-i\text{ times}}{\underbrace{\bm{0},\cdots,\bm{0}}})\cdot\bm{c}^{\top}=\bm{0}.$
@@
$(\underset{i-1\text{ times}}{\underbrace{\bm{0},\cdots,\bm{0}}},W_{i1},\underset{t-i+1\text{ times}}{\underbrace{\bm{0},\cdots,\bm{0}}})\cdot\bm{c}^{\top}=\bm{0}.$ (30)
@@
$\begin{split}&\begin{pmatrix}\bm{0}&\cdots&\bm{0}&W_{12}\\ N_{11}-N_{13}M_{13}^{-1}M_{11}&\cdots&N_{t1}-N_{13}M_{13}^{-1}M_{t1}&N_{12}-N_{13}M_{13}^{-1}M_{12}\end{pmatrix}\cdot\bm{c}^{\top}\\ &=\sum_{i=1}^{t}\begin{pmatrix}\bm{0}\\ N_{i1}-N_{13}M_{13}^{-1}M_{i1}\end{pmatrix}\cdot(\bm{c}_{i}|_{[s(r+\delta-1)]})^{\top}+\begin{pmatrix}W_{12}\\ N_{12}-N_{13}M_{13}^{-1}M_{12}\end{pmatrix}\cdot(\bm{c}^{*})^{\top}\\ &=-\sum_{i=1}^{t}\begin{pmatrix}W_{12}\\ N_{12}-N_{13}M_{13}^{-1}M_{12}\end{pmatrix}\cdot(\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]})^{\top}+\begin{pmatrix}W_{12}\\ N_{12}-N_{13}M_{13}^{-1}M_{12}\end{pmatrix}\cdot(\bm{c}^{*})^{\top}\\ &=\begin{pmatrix}W_{12}\\ N_{12}-N_{13}M_{13}^{-1}M_{12}\end{pmatrix}\cdot\left(\bm{c}^{*}-\sum_{i=1}^{t}\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]}\right)^{\top}\\ &=\bm{0},\end{split}$ (31)
where the second equation holds by $\overline{H}_{i}\cdot(\bm{c}_{i}|_{[(s+h)(r+\delta-1)]})^{\top}=\bm{0}$ for $i\in[t]$, and the last equation comes from (28). Combining (30) and (31), we have that $H_{F}\cdot\bm{c}^{\top}=\bm{0}$ for any $\bm{c}\in\mathcal{C}_{F}$. Therefore, $\mathcal{C}_{F}\subseteq\tilde{\mathcal{C}}$ and
[…]
Secondly, we claim that $\mathcal{C}_{F}$ is an optimal $(\ell+h)(r+\delta-1),\ell r,d]_{q}-$LRC. By the fact $\mathcal{C}_{F}\subseteq\tilde{\mathcal{C}}$, it is sufficient to show that $\mathcal{C}_{F}=\tilde{\mathcal{C}}$ and $\tilde{\mathcal{C}}$ is an optimal LRC. Note that $H_{F}$ is also the parity check matrix of the punctured code $\mathcal{C}|_{[(\ell+h)(r+\delta-1)]}$ from $\mathcal{C}$ by (26) and (27). Thus, $\tilde{\mathcal{C}}=\mathcal{C}|_{[(\ell+h)(r+\delta-1)]}$. Recall that $\mathcal{C}$ is an optimal LRC, which means
@@
On the other hand, $\tilde{\mathcal{C}}$ is an $(\ell+h)(r+\delta-1),\ell r,d(\tilde{\mathcal{C}})]_{a}-$LRC with $(r,\delta)_{a}$-locality, by the Singleton-type bound (3), we have
@@
Verify that the access cost is optimal: Combining (28) and (29), we derive that the write access cost is $h(r+\delta-1)$ and the read access cost is thrthr. According to Corollary 1, we have that
$\rho_{w}$ $\geqslant t\left(d+(t-1)k-1+\left(\left\lceil\frac{(t-1)k}{r}\right\rceil-1\right)(\delta-1)\right)-(t-1)n_{F}$
$=t(\delta+h(r+\delta-1)+(t-1)sr-1+(s(t-1)-1)(\delta-1))$
@@
$\Delta$ $=n_{F}-2d-(t-1)k+2-\left(\left\lceil\frac{(t-1)k}{r}\right\rceil-1\right)(\delta-1)$
$=(\ell+h)(r+\delta-1)-2(\delta+h(r+\delta-1))$
@@
$\rho_{r}$ $\geqslant t\left(sr-\Delta+(\delta-1)\left\lfloor\frac{\Delta}{r+\delta-1}\right\rfloor\right)$
$=t(sr+h(r+\delta-1)-s(r+\delta-1)+\delta-1+(\delta-1)(s-h-1))$
@@
The read and write access costs both achieve the lower bound in Corollary 1. Thus, the LRCC above is access-optimal, which completes the proof. ∎
In fact, the proof of Theorem 3 gives the following construction of access-optimal LRCCs.
Construction A:
Let $\mathcal{C}$ be an optimal $g(r+\delta-1),\ell r,d]_{q}-$LRC with $(r,\delta)_{a}$-locality, the parity check matrix $H$ of the from (1), where $\delta,g,\ell,r$ are positive integers with $\delta\geqslant 2$ and $g-\ell\geqslant 0$. Let $t$ be a positive integer such that $t\mid\ell$, $\ell/t>g$-$\ell$, and $h$ be a non-negative integer with $h\leqslant g$-$\ell$. The corresponding LRCC is given by the following steps:
  • First, generate initial optimal $(\ell/t+g-\ell)(r+\delta-1),\ell r/t]_{q}-$LRC $\mathcal{C}_{i}$ by means of the parity check matrix $H||_{\mathcal{P}_{i}\cup\mathcal{P}}$ as (25) for $i\in[t]$.
  • Second, establish the conversion procedure function as (28).
  • Finally, construct the final optimal $(\ell+h)(r+\delta-1),\ell r]_{q}-$LRC $\mathcal{C}_{F}$ with $(r,\delta)_{a}$-locality by means of the parity check matrix $H_{F}$ as (27).
@@
In general, the $t$ initial codes in Construction A could be the distinct codes. In the following case, the initial codes may be the same one. When $h=|\mathcal{P}|=g$-$\ell$ and $H||_{\mathcal{P}_{i}}=D_{i}\cdot H||_{\mathcal{P}_{1}}$ for $i\in[t]$ where $D_{i}$ is an invertible diagonal matrix formed as
$D_{i}=\begin{pmatrix}D_{i1}&\bm{0}_{s(\delta-1)\times hr}\\ \bm{0}_{hr\times s(\delta-1)}&D_{i2}\end{pmatrix}$.
In this case, both $H||_{\mathcal{P}_{i}\cup\mathcal{P}}=A_{i}\cdot H||_{\mathcal{P}_{1}\cup\mathcal{P}}$ and $H||_{\mathcal{P}_{1}\cup\mathcal{P}}$ can be the parity check matrices of initial code $\mathcal{C}_{i}$, where
$A_{i}=\begin{pmatrix}D_{i1}&\bm{0}_{s(\delta-1)\times h(\delta-1)}&\bm{0}_{s(\delta-1)\times hr}\\ \bm{0}_{h(\delta-1)\times s(\delta-1)}&I_{h(\delta-1)\times h(\delta-1)}&\bm{0}_{h(\delta-1)\times hr}\\ \bm{0}_{hr\times s(\delta-1)}&\bm{0}_{hr\times h(\delta-1)}&D_{i2}\end{pmatrix}$.
Therefore, for every $1\leqslant i\leqslant t$,
$\overline{H}_{i}=H||_{\mathcal{P}_{i}\cup\mathcal{P}}=\begin{pmatrix}W_{i1}&\bm{0}\\ \bm{0}&W_{12}\\ N_{i1}&N_{12}\end{pmatrix}$.
Since $\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{t}$ are the same code with parity check matrix $H||_{\mathcal{P}_{1}\cup\mathcal{P}}$, then modify the conversion procedure function (28) to
$(\bm{c}^{*})^{\top}=\begin{pmatrix}W_{12}\\ N_{12}\end{pmatrix}^{-1}\sum_{i=1}^{t}\left(\begin{pmatrix}I_{h(\delta-1)\times h(\delta-1)}&\bm{0}_{h(\delta-1)\times hr}\\ \bm{0}_{hr\times s(\delta-1)}&D_{i2}\end{pmatrix}\begin{pmatrix}W_{12}\\ N_{12}\end{pmatrix}\cdot(\bm{c}_{i}|_{[(s+h)(r+\delta-1)]\setminus[s(r+\delta-1)]})^{\top}\right),$ (35)
which shows (30) and (31) also hold. Moreover, the equations (30) and (31) hold in this case. Therefore, the new LRCC is access-optimal, and the initial and final codes are optimal LRCs, where the initial codes are equal.
V Explicit Constructions of LRCCs
In this section, we show two explicit constructions satisfying the final code preserves super-linear length or maximally repairable property when the initial codes have super-linear length or maximally repairable property respectively, based on Construction A by employing some base code $\mathcal{C}$ such that (1) $\mathcal{C}$ is an optimal $g(r+\delta-1),\ell r,d]_{q}-$LRC with $(r,\delta)_{a}$-locality, where $\delta\geqslant 2$ and $g-\ell\geqslant 0$; (2) $\mathcal{C}$ has the parity check matrix $H$ of the from (1).
@@ V-A Super-linear length
For MDS codes, the MDS conjecture implies that the length of MDS codes may not exceed $q+1$ ($q+2$ for $2\mid q$ and $k=3$), which means that for MDS convertible codes the length of the final codes may be upper bounded by $q+1$ ($q+2$ for $2\mid q$ and $k=3$). Analogously, Kong in [27] raises an interesting question about how large the code length, under the fixed size of the finite field, can be to ensure the existence of LRCCs with optimal access costs.
The trivial answer to this question is the length of the final code for LRCCs cannot exceed the optimal length of LRC referring to the known bounds in [6, 28].
Lemma 9 ([6, Theorem 2]):
Let $n=w(r+\delta-1)$, $\delta\geqslant 2$, $k=ur+v$, and additionally, $r|k$ or $u\geqslant 2(r+1-v)$, where all parameters are integers. Assume there exists an optimal $n,k,d]_{q}$ linear code $\mathcal{C}$ with all symbol $(r,\delta)$-locality, and define $t=\left\lfloor(d-1)/\delta\right\rfloor$. If $2t+1>4$, then
[…]
where wuw-u can also be rewritten as $w-u=\lfloor(d-1+v)/(r+\delta-1)\rfloor$.
In this subsection, we propose an explicit construction, which implies that the above bound is also tight asymptotically for LRCCs for some special cases where the final code has super-linear length when the initial codes have super-linear length.
Base Code A ([28, Construction A]): Let $d\geqslant\delta+1$ and $n=m(r+\delta-1)$, where $d,\delta,r,m$ are positive integers. For $i\in[m]$, let $G_{i}=\{g_{i,1},g_{i,2},\dots,g_{i,r+\delta-1}\}$ be a $(r+\delta-1)$-subset of $\mathbb{F}_{q}$. Then, for each $i\in[m]$, we can construct a $(d-1)\times(r+\delta-1)$ Vandermonde matrix with generating set $G_{i}$ of the form $(U_{i}^{\top},V_{i}^{\top})^{\top}$, where
@@
$H=\begin{pmatrix}U_{1}&\bm{0}&\cdots&\bm{0}\\ \bm{0}&U_{2}&\cdots&\bm{0}\\ \vdots&\vdots&\ddots&\vdots\\ \bm{0}&\bm{0}&\cdots&U_{m}\\ V_{1}&V_{2}&\cdots&V_{m}\end{pmatrix}$. (36)
@@
Lemma 10 ([28, Theorem III.4]):
Let $r+\delta\geqslant d\geqslant 2\delta+1$. Suppose that for any subset $S\subseteq[m]$ with $2\leqslant|S|\leqslant\lfloor\frac{d-1}{\delta}\rfloor$, then
@@
then any d1d-1 columns of $H$ given by (36) are linearly independent. As a result, the base code $\mathcal{C}$ is an optimal $n,k,d]_{q}-$LRC with $(r,\delta)_{a}$-locality.
@@
Recall that the access cost in Construction A also achieves the bound in Corollary 2. Therefore, we can derive the following lemma and theorem.
Lemma 11:
Let $m\geqslant v\geqslant 1$, and $i_{1}<i_{2}<\cdots<i_{v}$ where $i_{j}\in[m]$ for $j\in[v]$. Then the code $\mathcal{C}$ given by the parity check matrix $H||_{\{i_{1},i_{2},\dots,i_{v}\}}$ is an optimal $v(r+\delta-1),(v-1)r,r+\delta]_{q}-$LRC with $(r,\delta)_{a}$-locality, where $H$ is the parity check matrix of the code $\tilde{\mathcal{C}}$ from Base Code A satisfying $r+\delta=d\geqslant 2\delta+1$ and condition (37).
@@
Since $G_{i},i\in[m]$ satisfies condition (37), then $G_{j},j\in{\{i_{1},i_{2},\dots,i_{v}\}}\subseteq[m]$ satisfies condition (37). Also, the $\overline{d}$ for $\mathcal{C}$ satisfies $\overline{d}=d=r+\delta\geqslant 2\delta+1$. According to Lemma 10 and Lemma 3, the dimension $k(\mathcal{C})$ of $\mathcal{C}$ with parameters $v(r+\delta-1),k(\mathcal{C}),\overline{d}=r+\delta]_{q}$ satisfies
@@
Theorem 4:
Let $u$ be a positive integer. Let the code $\mathcal{C}$ be a Base Code A satisfying $r+\delta=d\geqslant 2\delta+1$ and condition (37). If $u\mid m-1$, then there is an access-optimal LRCC given by Construction A with parameters
@@
Moreover, the initial and final codes have super-linear length and both are optimal LRCs.
@@
In fact, $\mathcal{P}_{i}=\{(i-1)u+1,\dots,iu\}$ for $i\in[\frac{m-1}{u}]$, and $\mathcal{P}=\{m\}$. According to Corollary 2, the LRCC given by Construction A is access-optimal. According to Lemma 11, the initial codes are optimal $(u+1)(r+\delta-1),ur,r+\delta]_{q}-$LRC with $(r,\delta)_{a}$-locality.
It is well-known that the base code $\mathcal{C}$ can have super-linear length [28]. Therefore, the code length $m(r+\delta-1)$ satisfies the following:
@@
Since $(u+1)(r+\delta-1)=\frac{u+1}{m}\cdot m(r+\delta-1)$ and $\frac{u+1}{m}$ is a constant. Therefore, the initial codes can have super-linear length. ∎
@@
For the case $2\delta+1=d$, $r=\delta+1$, the parameters in Lemma 9 can be figured out as wu=u+1u=1w-u=u+1-u=1, $t=2$ and
@@
Thus, the code length $\Omega(q^{\delta})$ given by Theorem 4 is asymptotically optimal with respect to the bound in Lemma 9.
@@
Generally speaking, the $t$ initial codes in Theorem 4 could be the distinct codes. In the following case, the initial codes may be the same one. When arbitrary two of $\cup_{j\in\mathcal{P}_{i}}G_{j}$ for $i\in[t]$ belong to different cosets of one multiplicative group of $\mathbb{F}_{q}^{*}$ in Theorem 4, i.e.,